OpenStudy (anonymous):

Can anyone check my answer to a calc. 3 problem for me?

OpenStudy (anonymous):

Determine an equation of the line of intersection of the two planes: P1: 6x-3y+z=5 P2:+2x+2y+10z=10 My answer is: x = (-32t+40)/6 y = (-35+31t)/-3 z = t

OpenStudy (anonymous):

You solved it?

OpenStudy (anonymous):

Holy crap, how did you learn to input stuff into wolfram like that??